A youngs double slit experiment is conducted with slit seperation 10mm, where the screen is 2m away from the slits,if wavelength of light is used 6000A^0 if the wavelength is increased by1000A^0, and placed the whole apparatus in water of refractive index 4/3,the new fringe width in mm is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Frimge width is given as:
B = kD/d where k is the wavelength.
Now given,
B = 6000*2*10^-10/10.
Again apply formulae in water but the new wavelength would become:
k = 6000*(4/3)/1000